# Interconnection between density-regulation and stability in competitive ecological network

### v2

@@ 3.2 Analytical solution for local stability
We consider a random competitive community matrix $M$ with connectance $C$, the elements $M_{ij}$ are taken from half-negative distribution $-|X|$, where $E(\mathbf{X})=0$, $Var(\mathbf{X})=\sigma^{2}$. Then, $E(M_{ij})_{i\neq j}=-CE(|\mathbf{X}|)$, $Var(M_{ij})_{i\neq j}=C\sigma^{2}-C^{2}E^{2}(|\mathbf{X}|)$(Allesina and Tang, 2012). For simplify, we set $K_{i}=1$, $X_{i}^{*}=X^{*}$, $\theta_{i}=\theta_{1}(<1)$, $\theta_{2}(>1)$ and $M_{ii}=d_{1},d_{2}$ (for $\theta_{1}$ and $\theta_{2}$, respectively), with proportions $NP$ and $N(1-P)$, respectively. Note that, for the local stability criteria, the expected spectral abscissa ($\eta(M)$) should be negative.
@@ 4 Numerical simulations and results
Figure 1: Comparison between spectral abscissa (red line) and an average of the maximum eigenvalue (blue line) of $50$ community matrix as a measure of stability. The proportion of species ($p$) having density regulation ($\theta_{1}$) $0.1$ is plotted in the x-axis for panel (a). It is obvious that the density regulation ($\theta_{2}$) for complementary proportion ($1-p$) is $2$. The intrinsic growth rates for the two types of species are $5$ ($r_{1}$) and $0.25$ ($r_{2}$), respectively. The community size ($N$) and connectance ($C$) are $50$ and $1$, respectively. Competitive strength ($a_{ij}$) are chosen from $U[-0.2,0]$. The equal equilibrium size for all the species is $0.6$ ($X^{*}$). In panel (b), the proportion of species ($p$) having density regulation ($\theta_{1}$) $0.25$ is plotted on the x-axis. Density regulation ($\theta_{2}$) for complementary proportion ($1-p$) is $3$. The intrinsic growth rates for the two types of species are $3.6$ ($r_{1}$) and $0.3$ ($r_{2}$), respectively. The community size ($N$) and connectance ($C$) are $80$ and $0.9$, respectively. Competitive strength ($a_{ij}$) are chosen from $U[-0.1,0]$. The equal equilibrium size for all the species is $0.5$ ($X^{*}$).
@@ 4.1 Spectral distribution and Spectral abscissa of random community matrices
For an ensemble of randomly generated community matrices with fixed connectance($C$) and community size($N$), the expected maximum eigenvalue is given by $\eta(M)$ (expected spectral abscissa). For stability of the underlying topology, $\eta(M)$ should be negative. Figure 1(a) and 1(b) both are the same type; the only difference is that they are computed for different parameter sets. These figures exhibit how the maximum eigenvalue of a random competitive community matrix changes with the proportion $(p)$ of strong density regulated species (i.e., $\theta<1$). We investigate how $\eta(M)$ is regulated by connectance ($C$), community size ($N$), and the proportion $(p)$. $\eta(M)$ decreases as $p$ increases in the system for randomly generated competitive community matrices, depicted in Fig. 1. The spectral abscissa flips the sign from positive to negative for a sufficient proportion, making the system stable. We verify the result with numerical simulation (Fig. 1) by considering the average of maximum eigenvalues among fifty randomly generated community matrices for fixed connectance and community size. In Fig. 2, we illustrate how the proportion of two kinds of species ($p$) affects the eigenvalue distribution of random community structures. In the systems with strong density regulation at low or high population sizes ($p=0$ or $1$, respectively), the eigenvalue distribution appears to be elliptic. For $p$ lying between zero and one, eigenvalue distribution is subdivided into two ellipses, and the sizes of these ellipses are proportional to the proportion of weak and strong self-regulation at low population densities ($p$ and $1-p$). As the size of the ellipse close to zero increases, the length of the major axis increases, leading to eigenvalues bringing it close to zero and reducing its stability.
@@ Appendix A Derivation of the condition for stability
Let, the eigen values of $M$ be $\lambda_{i}(i=1,2,...,N)$. The expected spectral distribution of an ensemble of matrices can be written as $w(z)=E(\delta(z-\lambda_{i}))$ where $E(.)$ denotes expectation, and $\delta(.)$ denotes the Direct delta function.
Local stability of the equilibrium point $X^{*}$ depends on the spectral density abscissa $\eta(M)$ of $M.$ Here $\eta(M)=sup\{sup(Re(z):z\in support(w)\}$. The condition for the local stability of the system (1) in the main text is $\eta(M)$ should be less than zero.
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}}+\frac{1-P}{x^{2}}=1.$
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}}+\frac{1-P}{x^{2}}=1.$
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}}+\frac{1-P}{x^{2}}-1=0$
$\implies x^{2}P+(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}(1-P)-x^{2}(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}}))^{2}=0$
@@
$f(x)=x^{2}P+(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}(1-P)-x^{2}(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}}))^{2}$
we have $f(1)=P\left(1-(1+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}\right)<0$, and $f(0)=(1-P)(\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}>0$.
@@
$1=\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|X|))}})^{2}}+\frac{1-P}{x^{2}}.$

### v3

@@ 3.2 Analytical solution for local stability
We consider a random competitive community matrix $M$ with connectance $C$, the elements $M_{ij}$ are taken from half-negative distribution $-|\mathbf{X}|$, where $E(\mathbf{X})=0$, $Var(\mathbf{X})=\sigma^{2}$. Then, $E(M_{ij})_{i\neq j}=-CE(|\mathbf{X}|)$, $Var(M_{ij})_{i\neq j}=C\sigma^{2}-C^{2}E^{2}(|\mathbf{X}|)$(Allesina and Tang, 2012). For simplify, we set $K_{i}=1$, $X_{i}^{*}=X^{*}$, $\theta_{i}=\theta_{1}(<1)$, $\theta_{2}(>1)$ and $M_{ii}=d_{1},d_{2}$ (for $\theta_{1}$ and $\theta_{2}$, respectively), with proportions $NP$ and $N(1-P)$, respectively. Note that, for the local stability criteria, the expected spectral abscissa ($\eta(M)$) should be negative.
@@ 4 Numerical simulations and results
Figure 1: Comparison between spectral abscissa (red line) and an average of the real part of leading eigenvalue (blue line) of $50$ community matrix as a measure of stability. The proportion of species ($p$) having density regulation ($\theta_{1}$) $0.1$ is plotted in the x-axis for panel (a). It is obvious that the density regulation ($\theta_{2}$) for complementary proportion ($1-p$) is $2$. The intrinsic growth rates for the two types of species are $5$ ($r_{1}$) and $0.25$ ($r_{2}$), respectively. The community size ($N$) and connectance ($C$) are $50$ and $1$, respectively. Competitive strength ($a_{ij}$) are chosen from $U[-0.2,0]$. The equal equilibrium size for all the species is $0.6$ ($X^{*}$). In panel (b), the proportion of species ($p$) having density regulation ($\theta_{1}$) $0.25$ is plotted on the x-axis. Density regulation ($\theta_{2}$) for complementary proportion ($1-p$) is $3$. The intrinsic growth rates for the two types of species are $3.6$ ($r_{1}$) and $0.3$ ($r_{2}$), respectively. The community size ($N$) and connectance ($C$) are $80$ and $0.9$, respectively. Competitive strength ($a_{ij}$) are chosen from $U[-0.1,0]$. The equal equilibrium size for all the species is $0.5$ ($X^{*}$).
@@ 4.1 Spectral distribution and Spectral abscissa of random community matrices
For an ensemble of randomly generated community matrices with fixed connectance($C$) and community size($N$), the expected maximum real part of a eigenvalues is given by $\eta(M)$ (expected spectral abscissa). For stability of the underlying topology, $\eta(M)$ should be negative. Figure 1(a) and 1(b) both are the same type; the only difference is that they are computed for different parameter sets. These figures exhibit how the maximum real part of eigenvalues of a random competitive community matrix changes with the proportion $(p)$ of strong density regulated species (i.e., $\theta<1$). We investigate how $\eta(M)$ is regulated by connectance ($C$), community size ($N$), and the proportion $(p)$. $\eta(M)$ decreases as $p$ increases in the system for randomly generated competitive community matrices, depicted in Fig. 1. The spectral abscissa flips the sign from positive to negative for a sufficient proportion, making the system stable. We verify the result with numerical simulation (Fig. 1) by considering the average of maximum eigenvalues among fifty randomly generated community matrices for fixed connectance and community size. In Fig. 2, we illustrate how the proportion of two kinds of species ($p$) affects the eigenvalue distribution of random community structures. In the systems with strong density regulation at low or high population sizes ($p=0$ or $1$, respectively), the eigenvalue distribution appears to be elliptic. For $p$ lying between zero and one, eigenvalue distribution is subdivided into two ellipses, and the sizes of these ellipses are proportional to the proportion of weak and strong self-regulation at low population densities ($p$ and $1-p$). As the size of the ellipse close to zero increases, the length of the major axis increases, leading to eigenvalues bringing it close to zero and reducing its stability.
@@ Appendix A Derivation of the condition for stability
Let, the eigen values of $M$ be $\lambda_{i}(i=1,2,...,N)$. The expected spectral distribution of an ensemble of matrices can be written as $w(z)=E(\delta(z-\lambda_{i}))$ where $E(.)$ denotes expectation, and $\delta(.)$ denotes the Dirac delta function.
Local stability of the equilibrium point $X^{*}$ depends on the spectral density abscissa $\eta(M)$ of $M.$ Here $\eta(M)=sup\{Re(z):z\in support(w)\}$. The condition for the local stability of the system (1) in the main text is $\eta(M)$ should be less than zero.
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}}+\frac{1-P}{x^{2}}=1.$
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}}+\frac{1-P}{x^{2}}=1.$
@@
$\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}}+\frac{1-P}{x^{2}}-1=0$
$\implies x^{2}P+(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}(1-P)-x^{2}(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}}))^{2}=0$
@@
$f(x)=x^{2}P+(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}(1-P)-x^{2}(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}}))^{2}$
we have $f(1)=P\left(1-(1+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}\right)<0$, and $f(0)=(1-P)(\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}>0$.
@@
$1=\frac{P}{(x+\frac{\Delta}{\sqrt{N(C\sigma^{2}-C^{2}{E^{2}}(|\mathbf{X}|))}})^{2}}+\frac{1-P}{x^{2}}.$